What is the current in A, from a PV cell if dark saturation current, I0(Insat) =  1.000x10-10A, light generated current, IL=  0.500 A, ideality factor, n = 1, temperature, T = 300 K, and voltage, V =  0.500 V ? Express the result with 3 significant figures, include zero between the decimal point and do not write in the unit (it is already included in the question).
Answer #1

CONCEPT:

IN THE ABOVE FIGURES , 1ST FIGURE TELLS YOU ABOUT BASIC CIRCUIT OF ANY PV CELL HAVING A PHOTODIODE RECIEVING RAYS OF SUN AND POWERING UP THE LOAD AT ITS OUTPUT .

2ND FIGURE IS EXPANDED MODEL OF PV CELL HAVING A CONSTANT CURRENT SOURCE IL WHICH SIGNIFIES ABOUT LIGHT GENERATED CURRENT , ALSO A DIODE IS IN PARALLEL CARRYING ID DIODE CURRENT .

NOW , BY KCL AT NODE A ,

I= ( ID - IL )    // CURRENT FROM PV CELL(AT OUTPUT)

Where,    I= CURRENT DRAWN FROM PV CELL

   IL= PHOTOCURRENT OR LIGHT GENRATED CURRENT

ID = DIODE CURRENT = I0 [e (qV / nKb T) - 1)

where further, I0 = DARK SATURATION CURRENT

q = electric charge unit = 1.602 * 10-19 C

V = voltage

kb = boltzman constant = 1.381 * 10-23 J/K

T = temprature in Kelvin

n = ideality factor

NOW COMING TO QUESTION SOLUTION ,